Daniel has 8 wire panels that he will use the side of a new cage for his guinea pig, oscar. Each panal is 1 foot long
a) If he make the length of the base of the cage 3 feets and its total area will be 3 square feet,what will be width of the base of the cage be/ Show work
b) If he instated made all of side of the base 2 feet long, would Oscar have more room?.Shoe your work.

Answer by CubeyThePenguin(3113)

a) width = area/length = 3/3 = 1 ft

b) old area = (1)(3) = 3 ft^2
new area = (2)(2) = 4 ft^2

Yes, Oscar would have more room.